OpenStudy (anonymous):

Find the center and the radius of the circle: x2 - 6x + y2 + 1 = 0.

OpenStudy (ankit042):

(x-3)^2+y^2 +1-9=0

OpenStudy (ankit042):

so r = sqrt(8)

OpenStudy (anonymous):

how about to center though, thats what im having trouble with

OpenStudy (ankit042):

isn't it obvious now?

OpenStudy (ankit042):

(x-z)^2+(y-b)^2 = r^2 standard form where (a,b) is center you solve now